The filter cavity optics are in a drawer of B102. Please log and update when any of these optics are moved. Log changes in this log, and update the whereabouts on the wiki page:
https://wiki.mccullerlab.com/DCC/S2400001
These are tied to the DCC number as well (it will just link back, but the original spec was pulled from a LIGO DCC number):
https://dcc.ligo.org/LIGO-T2300191
Attached is an image where the optics are.
As you remove the existing cavity optics, the M1 of the cavity is a prior version of these. Please update also with the whereabouts of those 4 optics (They could be M0 or something).
The prior run that is T_1550=T_775=1% is in a box with the others now. It was in the same cabinet and easy to find. Be sure to return the one currently on the cavity to this box and document its history on the page.